OpenStudy (anonymous):

Line AB contains points A (3, −2) and B (1, 8). The slope of line AB is −5 -1/5 1/5 5

OpenStudy (helder_edwin):

given the points P(a,b) and Q(c,d) then the slope of the line PQ would be \[\large m=\frac{b-d}{a-c} \]

OpenStudy (anonymous):

@helder_edwin i dont understad?

OpenStudy (helder_edwin):

ok. see. u have the points \(A(\color{red}{3},\color{blue}{-2})\) and \(B(\color{green}{1},\color{orange}{8})\) so the slope is given by \[\large m=\frac{\color{blue}{-2}-\color{orange}{8}} {\color{red}{3}-\color{green}{1}}=\frac{-10}{2}=-5 \]

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) (3,-2) x1 = 3 and y1 = -2 (1,8) x2 = 1 and y2 = 8 now we sub slope = (8 - (-2) / (1 - 3) slope = (8 + 2) / -2 slope = 10/-2 slope = -5

OpenStudy (anonymous):

thank you both so much. @texaschic101 and @helder_edwin

OpenStudy (texaschic101):

sure thing :)